OpenStudy (anonymous):

x^2/a^2 + y^2/b^2 =1

OpenStudy (anonymous):

hmmmm whats the question?

OpenStudy (anonymous):

\(\dfrac{x^2}{a^2}+\dfrac{y^2}{b^2}=1\)